# RESCUE: Crowd Evacuation Simulation via Controlling SDM-United Characters

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: Our SDM-United Framework enables online simulation of personalized, physically realistic, and 3D-adaptive multi-agent evacuation scenarios.
@@ 1 Introduction
Evacuation simulations serve as valuable tools for assessing the likelihood of crowding and trampling incidents, estimating evacuation times, and supporting virtual reality escape training. However, no existing methods can simulate personalized and proxemics-aware 3D movements of hundreds of people online. This paper aims to develop an adaptive multi-agent 3D evacuation simulation framework. As shown in Fig. RESCUE: Crowd Evacuation Simulation via Controlling SDM-United Characters, our framework, called RESCUE (cRowd Evacuation Simulation via Controlling SDM-United charactErs), is capable of achieving online physically realistic simulations performing avoidance and personalized gait.
@@
Traditional crowd simulation methods [6, 22, 3] have explored various strategies in different scenarios. However, due to their simplified representations, these methods fail to integrate 3D motion with realistic behavior, resulting in physically implausible actions. Most crowd simulation methods cannot make decisions based on scene changes and are not designed for crowd evacuation simulation. Based on physics engines [18], character control [36, 26, 4, 29, 37] achieves autonomous 3D motion. However, these methods face the lack of personalization motions and thus result in falls and collisions in densely populated scenarios. Some motion generation methods [10, 31, 32, 43] based on diffusion models can generate diverse 3D motions, but they still face challenges in controllability and physical realism. Neither of these methods can directly produce realistic, personalized, and online evacuation motions. The detailed availabilities of the methods are summarized in Table 1.
@@
To achieve behavioral diversity, we introduce a Proxemics-aware Personalization Method with a personalized gait controller and optimized SFM to generate individualized behaviors for agents based on attributes such as ages and physical conditions. Additionally, we introduce Part-level Force Visualization, offering unprecedented insights into contact forces. Our experiments demonstrate that the proposed framework effectively adapts to uneven terrain, maintains balance during collisions and crowd congestion, and generates personalized behaviors. Although our experiments have focused on evacuation scenarios, the framework readily generalizes to a wide variety of multi agent environments, such as autonomous driving simulations in which dynamic vehicles and diverse pedestrians coexist. Our contributions can be summarized as follows:
@@
  • We propose a Personalized Gait Controller, enabling agents to generate personalized motions tailored to attributes such as age and physical condition. Additionally, a Part-level Forces Visualization is designed for analysis.
@@ 2.1 Crowd Simulation
Crowd simulation is a tool for analyzing and optimizing crowd dynamics. The aim is to capture key phenomena such as bottlenecks and panic-induced behaviors. Macroscopic models describe collective crowd movement using fluid dynamics [11, 5, 30, 8], while providing detailed insights through approaches. Social Force Models (SFM) [14, 15, 34] are among the most widely used methods, treating pedestrians as Newtonian particles influenced by the forces driving acceleration. Additional forces such as panic, sliding friction, and body compression have been introduced [6, 48]. However, these refinements often target specific scenarios. With advances in computational power, Agent-Based Models (ABM) have gained popularity [3, 7], simulating individuals as autonomous agents with unique decision-making capabilities. Although the above mentioned methods attempt to simulate the real-world escape, they rely on simplified 2D representations, where force feedback can only use unrealistic surrogate computations, leading to distorted speed outputs. Although several attempts at physically plausible 3D simulations have emerged [13, 46, 1, 9], these approaches still fall short in integrating personalized motion control and decision-making mechanisms. Most methods rely on generic movement patterns that overlook individual variations and fail to construct comprehensive strategies for handling crowd dynamics and escape scenarios.
@@ 2.3 Motion Generation
Motion generation aims to synthesize realistic and varied human motions. Recent advancements in diffusion models have significantly improved the diversity and quality of motion synthesis [31, 32]. Key research directions include text-driven motion generation, stylized motion synthesis [50, 2], scene-aware motion generation [40, 49, 23], and human-object interaction modeling [45]. Recent efforts [47, 33] have also integrated physical constraints to enhance the realism of generated motions. Trajectory-controlled [39, 43, 27] and multi-person motion synthesis [42] have been explored. However, challenges remain in generating long-duration, physically plausible motions, particularly in multiperson and contact-rich scenarios. Additionally, these methods usually take time to run and are not available online.
In conclusion, all existing methods fail to accurately simulate 3D multi-agent with personalized behaviors in densely populated escape scenarios. Therefore, we proposes the first SDM-unified evacuation simulation framework to generate realistic, personalized, and online 3D evacuation motions.
@@ 3.2 SDM-United Framework
The human brain follows a workflow of sensory, decision, and motor processes [28], and hence we develop our framework. Each person in evacuation process is abstracted as an agent that operates in an online loop, where the sensory module gathers observations, the decision module computes speeds, and the motor module executes motions, with updated states feeding back into the sensory module for continuous processing and making them available for next cycle.
@@
In 3D simulations, interaction dynamics differ from 2D point-based models due to physics engine integration. While 3D collision handling prevents interpenetration, human-to-human interactions require substantial adaptation—agents in 3D may stumble or fall under congestion that would merely slow movement in 2D models. Effective navigation requires detouring behavior rather than waiting or collision. To address these challenges, we introduce personalized SFM coefficient optimization and evasive forces specifically for 3D environments. Our decision process transforms agent $i$’s observations (position, speed, and nearby agents) into desired speed $\tilde{\boldsymbol{v}}_{i,t+1}\in\mathbb{R}^{3}$ for the next timestep.
@@
$\boldsymbol{F}_{\text{evasive}}=A\,\mathrm{sgn}\bigl{(}\boldsymbol{o}_{i}\!\cdot\!\boldsymbol{p}_{i}\bigr{)}\,\boldsymbol{p}_{i}\,,\vskip-6.99997pt$ (1)
@@
Personalized Optimization for SDM Coefficients. In crowd evacuation simulations, it seems not reasonable to use uniform social force coefficients, as individuals have varying escape abilities based on their attributes. Personalized coefficients better reflect these differences. Moreover, factors, for example friction in the physics engine, can cause individuals to move slower than the intended speed. To compensate for these, we adjust the target speed setting to achieve the desired speed in simulation. Specifically, $v_{\text{real}}$ is the empirically obtained real-world escape speed from literature (see Supplementary), $v_{\text{sim}}$ is the simulated speed, and $v_{\text{setting}}$ is the speed input to the engine. We run agents of different attributes on a straight path in the physics engine and optimize $v_{\text{setting}}$ so that $v_{\text{sim}}$ approximates $v_{\text{real}}$. We categorize agents into five groups, i.e. the young, middle-aged, elderly, patients, and disabled [^1] This classification is solely for scientific purposes, with no intention to offend or make assumptions based on age or disability.. Through this process, we determine personalized coefficients with an accuracy of 0.005 m/s for use in simulation. The algorithm details are presented in the Supplementary Materials.
@@ 3.4 Personalized Gait Control
Gait Frame Matching. Inspired by [21], a gait cycle can be fundamentally divided into 8 events [21]. We only use 4 events which are Initial Contact, Mid Stance, Opposite Initial Contact and Feet Adjacent. By calculating the distance between the two ankles, the image can be obtained similar to the sine-cosine function. The four gait events align with two consecutive cycles of the ankle distance waveform, corresponding to the peak, trough, peak, and trough, respectively. To ensure gait connectivity for gait frame matching, we assign gait values for all gait frames. We identify the keyframes of four events based on the above pattern and assign gait values of 0, 0.3, 0.5, and 0.75, while linearly interpolating the non-keyframes between them. We match personalized gait frames (excluding neutral style data in 100STYLE [19]) and non personalized gait frames (neutral style data in 100TYLE [19]) which have the same gait values as data pairs. The algorithm details are presented in the Supplementary Materials. When forming multiple candidate data frames, we select the two frames with the closest joint angles to match. During training, we also perform data augmentation by randomly assigning the same rotation to the root joints of the data pair.
Figure 3: Personalized Gait Converter
Network of Personalized Gait Converter. We use the CAMDM network [2] as backbone to implement the non-personalized action frame $\boldsymbol{a}_{i,t}$ to personalized action frame $\tilde{\boldsymbol{a}}_{i,t}$, see Figure 3. This network is a probability diffusion model. At each denoising step, the model takes as input a noisy motion sample $\boldsymbol{a}_{i,t}^{t}$, diffusion step $\mathbf{t}$, along with personalized gait label $c$, then learns to predict the original clean $\tilde{\boldsymbol{a}}_{i,t}^{0}$ (See Figure 3):
@@ 4.4 Ablation Study
Qualitative comparison. In qualitative comparison, we compare the pipeline of our framework after removing the 3D-adaptive SFM Decision Mechanism, the Personalizing Gait Control Motor, and both components together, which degrades into the PACER[29] baseline. For all methods, we first find waypoints. Figure 6 shows that, when it comes to excluding 3D adaptive SFM while keeping personalized gait control, the generated the agents fail to avoid each other, leading to collisions and crowd accumulation. As a result, some agents are unable to escape, and their escape speeds are nearly identical. This indicates that the 3DA-SFM Decision Mechanism effectively prevents potential collisions as much as possible and offers more rational escape processes tailored to individual attributes. Moreover, although removing the Personalized Gait Control Module allows most agents to fully escape, their gaits and movements become similar. While Personalized Gait Control Motor help to generate personalized motions for agents based on their attributes. In contrast, the complete pipeline can dynamically adjust escape routes based on evacuation speed tailored to each agent’s human attributes, minimizing collisions and ensuring personalized escape gaits that reflect individual characteristics. More details can be found in demo video and supplementary document.

### v2

@@ 1 Introduction
Evacuation simulations serve as valuable tools for assessing the likelihood of crowding and trampling incidents, estimating evacuation times, and supporting virtual reality escape training. However, no existing methods can simulate personalized and proxemics-aware 3D movements of hundreds of people online. This paper aims to develop an adaptive multi-agent 3D evacuation simulation framework. As shown in Figure LABEL:fig:fig_teaser, our framework, called RESCUE (cRowd Evacuation Simulation via Controlling SDM-United charactErs), is capable of achieving online physically realistic simulations performing avoidance and personalized gait.
@@
Traditional crowd simulation methods [6, 22, 3] have explored various strategies in different scenarios. However, due to their simplified representations, these methods fail to integrate 3D motion with realistic behavior, resulting in physically implausible actions. Most crowd simulation methods cannot make decisions based on scene changes and are not designed for crowd evacuation simulation. Based on physics engines [18], character control [36, 26, 4, 29, 37] achieves autonomous 3D motion. However, these methods face the lack of personalization motions and may result in falls and collisions in densely populated scenarios. Some motion generation methods [10, 31, 32, 43] based on diffusion models can generate diverse 3D motions, but they still face challenges in controllability and physical realism. Neither of these methods can directly produce realistic, personalized, and online evacuation motions. The detailed capabilities of the methods are summarized in Table 1.
@@
To achieve behavioral diversity, we introduce a Proxemics-aware Personalization Method with a personalized gait controller and an optimized SFM to generate individualized behaviors for agents based on attributes such as ages and physical conditions. Additionally, we introduce Part-level Force Visualization, offering unprecedented insights into contact forces. Our experiments demonstrate that the proposed framework effectively adapts to uneven terrain, maintains balance during collisions and crowd congestion, and generates personalized behaviors. Although our experiments have focused on evacuation scenarios, the framework readily generalizes to a wide variety of multi-agent environments, such as autonomous driving simulations in which dynamic vehicles and diverse pedestrians coexist. Our contributions can be summarized as follows:
@@
  • We propose a Personalized Gait Controller, enabling agents to generate personalized motions tailored to attributes such as age and physical condition. Additionally, a Part-level Force Visualization is designed for analysis.
@@ 2.1 Crowd Simulation
Crowd simulation is a tool for analyzing and optimizing crowd dynamics. The aim is to capture key phenomena such as bottlenecks and panic-induced behaviors. Macroscopic models describe collective crowd movement using fluid dynamics [11, 5, 30, 8], while providing detailed insights through approaches. Social Force Models (SFM) [14, 15, 34] are among the most widely used methods, treating pedestrians as Newtonian particles influenced by the forces driving acceleration. Additional forces such as panic, sliding friction, and body compression have been introduced [6, 48]. However, these refinements often target specific scenarios. With advances in computational power, Agent-Based Models (ABM) have gained popularity [3, 7], simulating individuals as autonomous agents with unique decision-making capabilities. Although the above-mentioned methods attempt to simulate real-world escape, they rely on simplified 2D representations, where force feedback can only use unrealistic surrogate computations, leading to distorted speed outputs. Although several attempts at physically plausible 3D simulations have emerged [13, 46, 1, 9], these approaches still fall short in integrating personalized motion control and decision-making mechanisms. Most methods rely on generic movement patterns that overlook individual variations and fail to construct comprehensive strategies for handling crowd dynamics and escape scenarios.
@@ 2.3 Motion Generation
Motion generation aims to synthesize realistic and varied human motions. Recent advances in diffusion models have significantly improved the diversity and quality of motion synthesis [31, 32]. Key research directions include text-driven motion generation, stylized motion synthesis [50, 2], scene-aware motion generation [40, 49, 23], and human-object interaction modeling [45]. Recent efforts [47, 33] have also integrated physical constraints to enhance the realism of generated motions. Trajectory-controlled [39, 43, 27] and multi-person motion synthesis [42] have been explored. However, challenges remain in generating long-duration, physically plausible motions, particularly in multiperson, contact-rich scenarios. In addition, these methods are usually time consuming and not suitable for online applications.
In conclusion, all existing methods fail to accurately simulate 3D multi-agent with personalized behaviors in densely populated escape scenarios. Therefore, we propose the first SDM-unified evacuation simulation framework to generate realistic, personalized, and online 3D evacuation motions.
@@ 3.2 SDM-United Framework
The human brain follows a workflow of sensory, decision, and motor processes [28], and hence we develop our framework. Each person in the evacuation process is abstracted as an agent that operates in an online loop, where the sensory module gathers observations, the decision module computes speeds, and the motor module executes motions, with updated states feeding back into the sensory module for continuous processing and making them available for next cycle.
@@
In 3D simulations, interaction dynamics differs from 2D point-based models due to physics engine integration. While 3D collision handling prevents interpenetration, human-to-human interactions require substantial adaptation—agents in 3D may stumble or fall under congestion that would merely slow movement in 2D models. Effective navigation requires detouring behavior rather than waiting or collision. To address these challenges, we introduce personalized SFM coefficient optimization and evasive forces specifically for 3D environments. Our decision process transforms agent $i$’s observations (position, speed, and nearby agents) into desired speed $\tilde{\boldsymbol{v}}_{i,t+1}\in\mathbb{R}^{3}$ for the next timestep.
@@
[…]
@@
Personalized Optimization for SDM Coefficients. In crowd evacuation simulations, it seems not reasonable to use uniform social force coefficients, as individuals have varying escape abilities based on their attributes. Personalized coefficients better reflect these differences. Moreover, factors, for example friction in the physics engine, can cause individuals to move more slowly than the intended speed. To compensate for these, we adjust the target speed setting to achieve the desired speed in simulation. Specifically, $v_{\text{real}}$ is the empirically obtained real-world escape speed from literature (see Supplementary), $v_{\text{sim}}$ is the simulated speed, and $v_{\text{setting}}$ is the speed input to the engine. We run agents of different attributes on a straight path in the physics engine and optimize $v_{\text{setting}}$ so that $v_{\text{sim}}$ approximates $v_{\text{real}}$. We categorize agents into five groups, i.e. the young, middle-aged, elderly, patients, and disabled [^1] This classification is solely for scientific purposes, with no intention to offend or make assumptions based on age or disability.. Through this process, we determine personalized coefficients with an accuracy of 0.005 m/s for use in simulation. The algorithm details are presented in the Supplementary Materials.
@@ 3.4 Personalized Gait Control
Gait Frame Matching. Inspired by [21], a gait cycle can be fundamentally divided into 8 events [21]. We only use 4 events which are Initial Contact, Mid Stance, Opposite Initial Contact and Feet Adjacent. By calculating the distance between the two ankles, the image can be obtained similar to the sine-cosine function. The four gait events align with two consecutive cycles of the ankle distance waveform, corresponding to the peak, trough, peak, and trough, respectively. To ensure gait connectivity for gait frame matching, we assign gait values for all gait frames. We identify the keyframes of four events based on the above pattern and assign gait values of 0, 0.3, 0.5, and 0.75, while linearly interpolating the non-keyframes between them. We match personalized gait frames (excluding neutral style data in 100STYLE [19]) and non-personalized gait frames (neutral style data in 100TYLE [19]) which have the same gait values as data pairs. When forming multiple candidate data frames, we select the two frames with the closest joint angles to match. The algorithm details are presented in the Supplementary Materials. During training, we also perform data augmentation by randomly assigning the same rotation to the root joints of the data pair.
Figure 3: Personalized Gait Converter.
Network of Personalized Gait Converter. We use the CAMDM network [2] as the backbone to implement the non-personalized action frame $\boldsymbol{a}_{i,t}$ to personalized action frame $\tilde{\boldsymbol{a}}_{i,t}$, see Figure 3. This network is a probability diffusion model. At each denoising step, the model takes as input a noisy motion sample $\boldsymbol{a}_{i,t}^{t}$, diffusion step $\mathbf{t}$, along with personalized gait label $c$, then learns to predict the original clean $\tilde{\boldsymbol{a}}_{i,t}^{0}$ (see Figure 3):
@@ 4.4 Ablation Study
Qualitative comparison. In qualitative comparison, we compare the pipeline of our framework after removing the 3D-adaptive SFM Decision Mechanism, the Personalizing Gait Control Motor, and both components together, which degrades into the PACER[29] baseline. For all methods, we first find waypoints. Figure 6 shows that, when it comes to excluding 3D adaptive SFM while keeping personalized gait control, the generated the agents fail to avoid each other, leading to collisions and crowd accumulation. As a result, some agents are unable to escape, and their escape speeds are nearly identical. This indicates that the 3DA-SFM Decision Mechanism effectively prevents potential collisions as much as possible and offers more rational escape processes tailored to individual attributes. Moreover, although removing the Personalized Gait Control Module allows most agents to fully escape, their gaits and movements become similar. Personalized Gait Control Motor helps to generate personalized motions for agents based on their attributes. In contrast, the complete pipeline can dynamically adjust escape routes based on evacuation speed tailored to each agent’s human attributes, minimizing collisions and ensuring personalized escape gaits that reflect individual characteristics. More details can be found in the demo video and supplementary document.